# Evaluating The Explainability of State-of-the-Art Machine Learning-based Online Network Intrusion Detection Systems

### v2

@@ IV-B Investigation Beyond Direct xAI Analysis
Under the systematic approach, for a given subset of data samples $D$, if $T$ is the set of features used in the selected root node-leaf node DT path decisions and $S$ is the set of features identified by kernel SHAP that contribute most to the classification decision for majority of the samples in $D$, then $T\subset S$ should hold true for the TRUSTEE DT explanation to agree with the SHAP explanation of the black-box model at the subset level. To form the set of features $S$, the mean absolute value of kernel SHAP values for multiple data samples are calculated for every feature, and subsequently the features are ranked according to their aggregate SHAP values to find the top features.
Inductive Bias: A number of ML models face the issue of underspecification. This term refers to the challenge of determining whether a model’s high performance is due to its genuine ability to capture essential patterns in the data or merely the result of certain inductive biases it has learned. These biases often appear as spurious correlations, a lack of ability to adapt to out-of-distribution (o.o.d.) data or shortcut learning strategies. The presence of such biases in trained models undermines their reliability and expected performance in real-world applications. Therefore, identifying these inductive biases is crucial for establishing trust in ML models. To detect the presence of an inductive bias in the ML-based NIDSs under consideration in this work, such as vulnerability to o.o.d. samples, we use tampered malicious traffic traces [10] which alter the attack segments of the original traffic traces by distributing the malicious packets in such a way that the packet rate remains within a specified limit.
@@ V Experimental Results
In this section, we present explanations of the four state-of-the-art ML-based online NIDSs (Kitsune, HorusEeye, ENIDrift, HELAD) generated using xAI methods such as TRUSTEE and kernel SHAP. For each pair of NIDS and xAI method, we wrote Python scripts for generating the explanation using the xAI method’s API and the modified source code of the NIDS (to make it compatible with the xAI method’s API implementation). Using the generated explanations, we present the most prominent features used by each NIDS model to make decisions on input data samples. We also investigate the relation between the explanations generated for a given NIDS by TRUSTEE and kernel SHAP by using the systematic approach for comparing TRUSTEE and SHAP explanations at the DT subset-level as explained in sub-section IV-B. Subsequently, we investigate the presence of inductive biases (e.g., vulnerability to o.o.d. data samples) in the NIDS implementations using tampered malicious traffic traces as explained in sub-section IV-B. Finally, we compare the explanations generated across the four NIDSs by TRUSTEE using the fidelity metric and by kernel SHAP using general xAI evaluation metrics.
Datasets: We use the Kitsune Mirai dataset [32] and the CICIDS-2017 dataset [23] for training/testing the ML-based NIDSs under consideration. The Mirai trace consists of $\sim$ 760k packets of synthetically generated attack in a network with nine IoT devices, first $\sim$120k packets consist of benign traffic, remaining $\sim$640k packets have anomalous traffic. The CICIDS-2017 dataset [23] consists of five days of packet capture (July 3 to July 7, 2017) from a small-scale enterprise-like network, with the first day consisting of benign background traffic and the other days consisting of attack traffic (brute force SSH, DoS, Heartbleed, web attack, infiltration, etc.). It consists of significant drift over the data collection period in terms of different attack tools used. Below, we present the results using Mirai dataset only. Since the results obtained with CICIDS-2017 dataset are similar to those obtained with the Mirai dataset, we include them in the Appendix Section A.
@@ V-A Case Study- Kitsune $\times$ Mirai
The top three prominent features that Kitsune uses to determine an anomaly are:
@@
TRUSTEE vs SHAP Analysis: Using the systematic approach for comparing TRUSTEE and SHAP explanations as explained in sub-section IV-B, since the set of features used in the benign root node-leaf node path in the TRUSTEE DT is not a subset of the set of top contributing features computed by kernel SHAP summary analysis ($HH\_0.1\_weight\_0$ is the only common feature between the two sets), TRUSTEE AND SHAP explanations do not agree with each other at the subset level for Kitsune NIDS.
@@
Detection of Inductive Bias: It was shown in [10] that Kitsune NIDS is vulnerable to o.o.d. samples when used with the Mirai dataset introduced in [32]. The authors evaluated the following threshold ranges (in packets per second) for ARP requests originating from devices infected with Mirai: (1) 10-50, (2) 30-70 and (3) 50-90. The original trace as well as the tampered traces are shown in Fig. 4. It can be observed that tampering had a significant effect on the difference between benign and attack traffic in terms of number of packets per second. Kitsune was executed on each trace with an identical number of training samples. The results showed that the RMSE values generated by Kitsune NIDS decreased significantly with a drop in the attack traffic volume, reaching the values corresponding to benign traffic.
@@ V-B Case Study- HorusEye $\times$ Mirai
SHAP Analysis: Using kernel SHAP analysis, Fig. 6 shows a summary of how the top features in Mirai dataset impact the HorusEye NIDS model’s output using a beeswarm plot for randomly selected 1000 data samples corresponding to the benign root node-leaf node path in the TRUSTEE DT (Fig. 5). Features such as $HH\_0.1\_pcc\_0\_1$, $HH\_1\_mean\_0$ and $HH\_0.01\_std\_0$ have the most impact on the prediction value, in that order. $HH\_0.1\_pcc\_0\_1$ corresponds to the correlation coefficient between two packet size streams aggregated by traffic sent between a set of source and destination IP addresses with time window 0.1 (10 seconds), $HH\_1\_mean\_0$ corresponds to the mean of packet sizes aggregated by traffic sent between a set of source and destination IP addresses with time window 1 (1.5 seconds) and $HH\_0.01\_std\_0$ corresponds to the standard deviation of packet sizes aggregated by the traffic sent between a set of source and destination IP addresses with time window 0.01 (1 minute).
@@
TRUSTEE vs SHAP Analysis: Using the systematic approach for comparing TRUSTEE and SHAP explanations as explained in sub-section IV-B, the set of features used in the benign root node-leaf node path in the TRUSTEE DT is a subset of the set of top contributing features computed by kernel SHAP summary analysis. Therefore, TRUSTEE AND SHAP explanations agree with each other at the subset level for HorusEye NIDS.
Detection of Inductive Bias: Following a similar approach as the Kitsune case, three separate tampered traces with different attack traffic volumes were obtained from the Mirai dataset and HorusEye was executed on each trace. In Fig. 7(a), we show the results for first 200,000 packets of each traces. In Fig. 7(b), we also compare the expected RMSE (produced by HorusEye in the original trace) and the predicted RMSE for each tampered trace. The results show that the RMSE values generated by HorusEye decrease with a drop in the attack traffic volume. However, the difference between the RMSE values produced by the original and tampered traces is not as pronounced as in the case of Kitsune NIDS. This experiment demonstrates that the Mirai use case from HorusEye is vulnerable to o.o.d. samples, though to a lesser extent than Kitsune. It should be noted here that HorusEye [33] authors had recognized this issue (without an empirical analysis though) by stating that the frequency of Mirai attacks is too high in the original dataset to reflect real botnet infection. Therefore, the authors had randomly down-sampled the Mirai attack traffic in Kitsune dataset to consist of 40,000 packets during HorusEye’s performance evaluation.
@@ V-C Case Study- ENIDrift $\times$ Mirai
Detection of Inductive Bias: Following a similar approach as the Kitsune case, three separate tampered traces with different attack traffic volumes were obtained from the Mirai dataset and ENIDrift was executed on each trace. In Fig. 10(a), we show the results for first 120,000 packets of each traces. In Fig. 10(b), we also compare the expected (weighted) ensemble scores (produced by ENIDrift in the original trace) and the predicted (weighted) ensemble scores for each tampered trace. The results show that the ensemble scores produced by ENIDrift do not have a discernable pattern of dependence on the volume of the attack traffic encountered. Since the distance of the dots for tampered traces from the diagonal line in Fig. 12(b) does not follow any pattern, the effect that tampering with the Mirai attack traces had on the predicted ensemble scores and hence, whether the Mirai use case from ENIDrift is vulnerable to o.o.d. samples or not, remains inconclusive.
@@ V-D Case Study- HELAD $\times$ Mirai
Detection of Inductive Bias: Following a similar approach as the Kitsune case, three separate tampered traces with different attack traffic volumes were obtained from the Mirai dataset and HELAD was executed on each trace. In Fig. 12(a), we show the results for first 200,000 packets of each traces. In Fig. 12(b), we also compare the expected anomaly scores (produced by HELAD in the original trace) and the predicted anomaly scores for each tampered trace. The results show that the anomaly scores produced by HELAD for the original and tampered traces overlap with each other and therefore, the scores do not have a significant dependence on the volume of the attack traffic encountered. Since the dots for tampered traces are close to the diagonal line in Fig. 12(b), tampering with the Mirai attack traces had limited outcome on the predicted anomaly scores. This experiment demonstrates that the Mirai use case from HELAD is not vulnerable to o.o.d. samples to a significant extent.
@@ V-E Comparative Analysis
In this sub-section, we use xAI evaluation metrics to quantify the effectiveness of the explanations generated for the ML-based NIDSs under consideration by TRUSTEE and kernel SHAP. We use those metrics to compare the explanations across NIDSs. Currently, the existing implementations of xAI evaluation metrics support kernel SHAP only and none of the implementations support TRUSTEE. TRUSTEE’s implementation comes with its own metric called fidelity which quantifies how accurate the TRUSTEE DT is with respect to the original black-box ML model.
@@
Using the CICIDS-2017 dataset, if we compare the fidelity values obtained for the NIDSs considered (refer Appendix Section A), without any pruning, the highest fidelity values are obtained for ENIDrift (0.929) out of the four NIDSs. With top-k pruning ($k=10$), the highest fidelity values are obtained for Kitsune (0.919) out of the four NIDSs. Since HorusEye and HELAD NIDSs yield negative and zero fidelity values, TRUSTEE DT explanations do not serve much purpose. With top-k pruning, TRUSTEE is most useful for explaining the decisions of Kitsune NIDS followed by ENIDrift NIDS. Across the datasets, with top-k pruning, Kitsune NIDS yields the highest fidelity values. The comparison of fidelity values obtained during TRUSTEE analysis of the ML-based NIDSs considered in this paper is shown in Table V.
Online NIDS Fidelity (Without pruning) Fidelity (with top-k pruning, k=10)
@@
TABLE V: Comparison of TRUSTEE fidelity values with and without pruning for ML-based online NIDSs considered
SHAP evaluation metrics: We experimented with multiple implementations of xAI evaluation metrics (e.g., faithfulness, stability, monotonicity) to compare the kernel SHAP explanations across the four NIDSs under consideration in our paper. Some of those implementations are OpenXAI  [^1] https://github.com/AI4LIFE-GROUP/OpenXAI, XAI-bench  [^2] https://github.com/abacusai/xai-bench and Quantus  [^3] https://github.com/understandable-machine-intelligence-lab/Quantus. However, after extensive experimentation with those implementations, we found out that none of them support custom ML models such as those used for NIDSs. They support out-of-the-box ML models such as linear regression, MLP and ANN only.

### v3

@@ IV-B Investigation Beyond Direct xAI Analysis
Under the systematic approach, for a given subset of data samples $D$, if $T$ is the set of features used in the selected root node-leaf node DT path decisions and $S$ is the set of features identified by kernel SHAP that contribute most to the classification decision for majority of the samples in $D$, then $T\subset S$ should hold true for the TRUSTEE DT explanation to agree with the SHAP explanation of the black-box model at the subset level. To form the set of features $S$, the mean absolute value of kernel SHAP values for multiple data samples are calculated for every feature, and subsequently the features are ranked according to their aggregate SHAP values to find the top features. For a black-box NIDS model, the local explanation (SHAP) may agree with the global explanation (TRUSTEE) depending on the data subset on which they are executed. However, not all local explanations may agree with the global explanation. To evaluate the level of agreement between TRUSTEE and SHAP for an NIDS trained on a given dataset, we propose the following approach:
  1. 1.
    Sample different subsets from the original evaluation dataset by selecting different root node-leaf node paths in the TRUSTEE DT.
  2. 2.
    Use each sampled data subset for executing kernel SHAP on the trained NIDS.
  3. 3.
    We propose the TRUSTEE-SHAP agreement score on a data subset $D$ which can be calculated as:
    $\alpha_{TRUSTEE-SHAP}^{D}=\frac{n(T\subset S)}{n(T)},$ (3)
    where $n(T\subset S)$ refers to the number of TRUSTEE DT features which are a subset of SHAP features and $n(T)$ is the total number of TRUSTEE DT features corresponding to the selected data subset $D$.
We also define the m-point average TRUSTEE-SHAP agreement score for an NIDS as the average of TRUSTEE-SHAP agreement scores for $m$ different data subsets. It can be expressed as:
$A_{TRUSTEE-SHAP}^{m}=\frac{1}{m}\sum_{m}\alpha_{TRUSTEE-SHAP}^{D_{m}}$ (4)
Inductive Bias: A number of ML/DL models face the issue of underspecification. This term refers to the challenge of determining whether a model’s high performance is due to its genuine ability to capture essential patterns in the data or merely the result of certain inductive biases it has learned. These biases often appear as spurious correlations, a lack of ability to adapt to data samples with variable statistics, shortcut learning strategies, etc. The presence of such biases in trained models undermines their reliability and expected performance in real-world applications. Therefore, identifying these inductive biases is crucial for establishing trust in ML/DL models. To detect the presence of an inductive bias in the DL-based NIDSs under consideration in this work, such as vulnerability to variable packet rate samples, we use tampered malicious traffic traces [7] which alter the attack segments of the original traffic traces by distributing the malicious packets in such a way that the packet rate remains within a specified limit.
@@ V Experimental Results
In this section, we present explanations of state-of-the-art DL-based NIDSs generated using xAI methods such as TRUSTEE and kernel SHAP. For each pair of NIDS and xAI method, we wrote Python scripts for generating the explanation using the xAI method’s API and the modified source code of the NIDS (to make it compatible with the xAI method’s API implementation). Using the generated explanations, we present the most prominent features used by each NIDS model to make decisions on input data samples. We also investigate the relation between the explanations generated for a given NIDS by TRUSTEE and kernel SHAP by using the systematic approach for comparing TRUSTEE and SHAP explanations at the DT subset-level as explained in sub-section IV-B. Subsequently, we investigate the presence of inductive biases (e.g., vulnerability to variable packet rate samples) in the NIDS implementations using tampered malicious traffic traces as explained in sub-section IV-B. Finally, we compare the explanations generated across the four NIDSs by TRUSTEE using the fidelity metric and by kernel SHAP using general xAI evaluation metrics.
NIDSs Evaluated: We have used all the state-of-the-art DL-based NIDSs presented in Section III-C (Kitsune, HorusEeye, ENIDrift and HELAD) for our analysis. When we refer to HorusEye from this point onwards, we are referring to its Magnifier component which performs deep learning-based anomaly detection.
Datasets: We use the Kitsune Mirai dataset [5] and the CICIDS-2017 dataset [17] for training/testing the DL-based NIDSs under consideration. The Mirai PCAP trace consists of $\sim$ 760k packets of synthetically generated attack in a network with nine IoT devices, first $\sim$120k packets consist of benign traffic, remaining $\sim$640k packets have anomalous traffic. We use the Kitsune code to extract features from the PCAP trace through the DIS method. The CICIDS-2017 dataset [17] consists of five days of packet capture (July 3 to July 7, 2017) from a small-scale enterprise-like network, with the first day consisting of benign background traffic and the other days consisting of attack traffic (brute force SSH, DoS, Heartbleed, web attack, infiltration, etc.). It consists of significant drift over the data collection period in terms of different attack tools used. The CICIDS-2017 dataset provides PCAP traces, labeled network traffic flows and CSV files consisting of pre-selected features extracted from traffic flows. We use the raw PCAP traces directly and extract features from them using the Kitsune DIS method to maintain consistency. Below, we present the results using Mirai dataset only. Since the results obtained with CICIDS-2017 dataset are similar to those obtained with the Mirai dataset, we include them in the Appendix Section A.
@@ V-A Case Study- Kitsune $\times$ Mirai
The fidelity score presents a way for security practitioners to gauge the effectiveness of the explanation generated for an NIDS. A security practitioner prioritising just explainability should select the NIDS with the highest fidelity score. However, a typical practitioner interested in deploying a DL-based NIDS will prioritise not just its explainability but also its its attack detection performance. Traditionally, the detection performance of ML/DL-based NIDSs has been measured in terms of metrics such as accuracy, precision, recall and F1 score. An example of the metric a security practitioner can use while selecting a DL-based NIDS is a weighted combination of the NIDS’s average F1 score for different kinds of attacks and TRUSTEE fidelity, expressed as below:
$S=\alpha*(F1-score)+(1-\alpha)*(Fidelity),0\leq\alpha\leq 1$ (5)
By varying the value of the weight, $\alpha$, the security practitioner can prioritise between detection performance and fidelity.
Based on the TRUSTEE analysis results obtained earlier, the top three prominent features that Kitsune uses to determine an anomaly are:
@@
Information about the top prominent packet-based features used by a DL-based NIDS for anomaly detection makes it easier for a security practitioner interested in deploying the NIDS to trust their internal decision making process. It also enables practitioners to compare NIDSs based on their respective top features. However, practitioners need to select only those NIDSs for comparison whose fidelity is above a certain minimum threshold. If the fidelity falls below that threshold, it means that the DT explanation generated for the corresponding NIDS is not reliable.
@@
Detection of Inductive Bias: It was shown in [7] that Kitsune NIDS is vulnerable to variable packet rate samples when used with the Mirai dataset introduced in [5]. The authors evaluated the following threshold ranges (in packets per second) for ARP requests originating from devices infected with Mirai: (1) 10-50, (2) 30-70 and (3) 50-90. The original trace as well as the tampered traces are shown in Fig. 4. It can be observed that tampering had a significant effect on the difference between benign and attack traffic in terms of number of packets per second. Kitsune was executed on each trace with an identical number of training samples. The results showed that the RMSE values generated by Kitsune NIDS decreased significantly with a drop in the attack traffic volume, reaching the values corresponding to benign traffic.
@@ V-B Case Study- HorusEye $\times$ Mirai
SHAP Analysis: Using kernel SHAP analysis, Fig. 6 shows a summary of how the top features in Mirai dataset impact the HorusEye NIDS model’s output using a beeswarm plot for randomly selected 1000 data samples corresponding to the benign root node-leaf node path in the TRUSTEE DT (Fig. 5). Features such as $HH\_1\_std\_0$, $HH\_5\_std\_0$ and $HpHp\_0.01\_magnitude\_0\_1$ have the most impact on the prediction value, in that order. $HH\_1\_std\_0$ and $HH\_5\_std\_0$ correspond to the standard deviation of packet sizes aggregated by the traffic sent between a set of source and destination IP addresses with time windows 1 (1.5 seconds) and 5 (100 ms) respectively, while $HpHp\_0.01\_magnitude\_0\_1$ corresponds to the absolute magnitude of two packet size streams aggregated by the traffic sent between a set of source and destination IP addresses with time window 0.01 (1 minute).
@@
Detection of Inductive Bias: Following a similar approach as the Kitsune case, three separate tampered traces with different attack traffic volumes were obtained from the Mirai dataset and HorusEye was executed on each trace. In Fig. 7(a), we show the results for first 200,000 packets of each traces. In Fig. 7(b), we also compare the expected RMSE (produced by HorusEye in the original trace) and the predicted RMSE for each tampered trace. The results show that the RMSE values generated by HorusEye decrease with a drop in the attack traffic volume. However, the difference between the RMSE values produced by the original and tampered traces is not as pronounced as in the case of Kitsune NIDS. This experiment demonstrates that the Mirai use case from HorusEye is vulnerable to variable packet rate samples, though to a lesser extent than Kitsune. It should be noted here that HorusEye [28] authors had recognized this issue (without an empirical analysis though) by stating that the frequency of Mirai attacks is too high in the original dataset to reflect real botnet infection. Therefore, the authors had randomly down-sampled the Mirai attack traffic in Kitsune dataset to consist of 40,000 packets during HorusEye’s performance evaluation.
@@ V-C Case Study- ENIDrift $\times$ Mirai
Detection of Inductive Bias: Following a similar approach as the Kitsune case, three separate tampered traces with different attack traffic volumes were obtained from the Mirai dataset and ENIDrift was executed on each trace. In Fig. 10(a), we show the results for first 120,000 packets of each traces. In Fig. 10(b), we also compare the expected (weighted) ensemble scores (produced by ENIDrift in the original trace) and the predicted (weighted) ensemble scores for each tampered trace. The results show that the ensemble scores produced by ENIDrift do not have a discernable pattern of dependence on the volume of the attack traffic encountered. Since the distance of the dots for tampered traces from the diagonal line in Fig. 12(b) does not follow any pattern, the effect that tampering with the Mirai attack traces had on the predicted ensemble scores and hence, whether the Mirai use case from ENIDrift is vulnerable to variable packet rate samples or not, remains inconclusive.
@@ V-D Case Study- HELAD $\times$ Mirai
Detection of Inductive Bias: Following a similar approach as the Kitsune case, three separate tampered traces with different attack traffic volumes were obtained from the Mirai dataset and HELAD was executed on each trace. In Fig. 12(a), we show the results for first 200,000 packets of each traces. In Fig. 12(b), we also compare the expected anomaly scores (produced by HELAD in the original trace) and the predicted anomaly scores for each tampered trace. The results show that the anomaly scores produced by HELAD for the original and tampered traces overlap with each other and therefore, the scores do not have a significant dependence on the volume of the attack traffic encountered. Since the dots for tampered traces are close to the diagonal line in Fig. 12(b), tampering with the Mirai attack traces had limited outcome on the predicted anomaly scores. This experiment demonstrates that the Mirai use case from HELAD is not vulnerable to variable packet rate samples to a significant extent.
@@ V-E Comparative Analysis
In this sub-section, we use xAI evaluation metrics to quantify the effectiveness of the explanations generated for the DL-based NIDSs under consideration by TRUSTEE and kernel SHAP. We use those metrics to compare the explanations across NIDSs. Currently, the existing implementations of objective xAI evaluation metrics as explained in Section III-B support kernel SHAP only and none of the implementations support TRUSTEE. TRUSTEE’s implementation comes with its own metric called fidelity which quantifies how accurate the TRUSTEE DT is with respect to the original black-box ML model. We also use the TRUSTEE-SHAP agreement scores proposed in Section IV-B to compare the level of agreement between TRUSTEE DT and kernel SHAP explanations generated for the DL-based NIDSs considered.
@@
Using the CICIDS-2017 dataset, if we compare the fidelity values obtained for the NIDSs considered (refer Appendix Section A), without any pruning, the highest fidelity values are obtained for ENIDrift (0.929) out of the four NIDSs. With top-k pruning ($k=10$), the highest fidelity values are obtained for Kitsune (0.919) out of the four NIDSs. Since HorusEye and HELAD NIDSs yield negative and zero fidelity values, TRUSTEE DT explanations do not serve much purpose. With top-k pruning, TRUSTEE is most useful for explaining the decisions of Kitsune NIDS followed by ENIDrift NIDS. Across the datasets, with top-k pruning, Kitsune NIDS yields the highest fidelity values. The comparison of fidelity values obtained during TRUSTEE analysis of the DL-based NIDSs considered in this paper is shown in Table V.
NIDS Name Fidelity (Without pruning) Fidelity (with top-k pruning, k=10)
@@
TABLE V: Comparison of TRUSTEE fidelity values with and without pruning for DL-based NIDSs considered
SHAP evaluation metrics: We experimented with multiple implementations of objective xAI evaluation metrics (e.g., completeness, faithfulness, stability, monotonicity) to compare the kernel SHAP explanations across the four NIDSs evaluated in our experiments. Some of those implementations are OpenXAI  [^1] https://github.com/AI4LIFE-GROUP/OpenXAI, XAI-bench  [^2] https://github.com/abacusai/xai-bench and Quantus  [^3] https://github.com/understandable-machine-intelligence-lab/Quantus. However, after extensive experimentation with those implementations, we found out that none of them support custom ML models such as those used for NIDSs. They support out-of-the-box ML models such as linear regression, MLP and ANN only.
Level of TRUSTEE-SHAP Agreement: We set the value of $m=3$ for the $m$-point average TRUSTEE-SHAP agreement score. Each of the data subsets used for calculation of TRUSTEE-SHAP agreement scores were required to have a minimum number of 300 samples. The results are shown in Table VI. HorusEye, with a score of $0.381$, exhibits the highest level of TRUSTEE-SHAP agreement, followed by Kitsune, with a score of $0.167$ while ENIDrift has a score of $0.0$. We also find that the TRUSTEE-SHAP agreement scores obtained for an NIDS for the different data subsets vary significantly. For example, for Kitsune, the scores obtained for three different data subsets were $0.4$, $0.6$ and $0.1$. This supports our earlier assumption that the level of agreement between TRUSTEE DT explanation and SHAP explanation for a blackbox NIDS model may depend on the specific data subset which is used to conduct the TRUSTEE-SHAP comparison.
NIDS $m$-point avg. TRUSTEE-SHAP agreement score (m=3)
Kitsune 0.167
HorusEye 0.381
ENIDrift 0.0
HELAD n/a
TABLE VI: Comparison of $m$-point average TRUSTEE-SHAP agreement scores for DL-based SOTA NIDSs trained with Kitsune Mirai dataset